OpenStudy (anonymous):

evaluate: integral from 3 to infinity: 1/(x-2)^(3/2) dx

OpenStudy (anonymous):

\[\int\limits_{3}^{\infty?} (1)\div ((x-2)^{(3/2)}) dx\]

OpenStudy (anonymous):

\[\lim_{n \rightarrow \inf} \int\limits_{3}^{n}(1/((n-2)^(3/2))dn\] = \[\lim_{n \rightarrow \inf} [-2/(\sqrt{n-2}) - (-2/\sqrt{3-2} )\] =0+2=2

OpenStudy (anonymous):

I skipped all of the integral steps and went straight to evaluating them... let me know if you have questions on the in between steps.

OpenStudy (anonymous):

ah yes! thats the part i got stuckk! thank u :)

OpenStudy (anonymous):

so you got it?

OpenStudy (anonymous):

where is the -2 in the numerator coming from?

OpenStudy (anonymous):

\[\int_3^\infty (x-2)^{-3/2} dx=\frac{(x-2)^{-1/2}}{-1/2}+C\] \[(x-2)^{-1/2}=\frac{1}{\sqrt{x-2}}\], which approaches 0 as x approaches infinity. Hence the integral is \[0-\frac{(3-2)^{-1/2}}{-1/2}=2\]

OpenStudy (anonymous):

I used u substitution for the integral and integral of u^(-3/2) = -2/sqrt(u)